# Exergetic Port-Hamiltonian Systems Modeling Language

### v1

@@ 8 Thermodynamic consistency
As already touched upon in the introduction, the GENERIC establishes a general structure, also called metriplectic structure, for combining reversible and irreversible dynamics in a way that respects fundamental principles from thermodynamics, namely the first law, the second law, and Onsager reciprocal relations.
The vector field which determines the dynamics of a metriplectic system is given as the sum of a reversible Hamiltonian vector field that conserves energy due to the skew-symmetry of the Poisson bivector and a vector field of an irreversible gradient flow that has non-negative entropy production due to the non-negative definiteness of the metric tensor. Further, so-called non-interacting or degeneracy conditions ensure that the reversible dynamics conserves entropy and the irreversible dynamics conserves energy. The first law then holds since both parts conserve energy and the second law holds because the irreversible part has non-negative entropy production, while the reversible part conserves entropy, see e.g. Morrison (1984); Grmela and Öttinger (1997). Onsager reciprocal relations hold, as the symmetry of the metric tensor implies equal cross-effect coefficients, see e.g. Mielke et al. (2016).
Usually, the reversible part and the irreversible part both satisfy further degeneracy conditions that ensure the conservation of further quantities such as volume or mass. Given that (some of) these quantities, together with entropy, are present in an exergy reference environment that is adjoined to the metriplectic system, it follows that the reversible dynamics conserve exergy and the irreversible dynamics are characterized by a non-negative exergy destruction. This immediately follows from the definition of the exergy storage function analogous to 6.2.
The argument for EPHS is very similar in nature, although it is in a sense reversed. To keep it short and simple, we focus on isolated systems.
@@
The dynamics of any isolated EPHS respects the first law, the second law, and Onsager reciprocal relations. ∎
@@
We argue in terms of components and their interconnection.
@@
    According to 6.9, all reversible components conserve exergy as well as all quantities of the reference environment.
    Considering the definition of exergy in 6.2, it follows immediately that all reversible components also conserve energy.
@@
    According to 6.14, all irreversible components are defined by a non-negative exergy destruction rate and they conserve energy as well as all quantities of the reference environment, except for entropy.
    Considering the definition of exergy, it follows that all irreversible components also have a non-negative entropy production rate.
@@
    According to 4.2, any interconnection of the components conserves exergy and all quantities (of the reference environment), since flows balance at equal efforts.
    Considering the definition of exergy, it follows that any interconnection of the components also conserves energy.
The first law holds since all three aspects conserve energy. The second law holds because the irreversible components have a non-negative entropy production rate, while the reversible components and the interconnection conserve entropy. The reciprocal relations hold because, according to 6.12, thermodynamic fluxes (flows) depend on thermodynamic forces (efforts) via a symmetric locally-linear operator. ∎
We note that the argument is reversed because, according to Definitions 6.4 and 7.2, the semantics of EPHS is defined in terms of a storage function that is interpreted as the exergy of the system, as opposed to two separate functions representing the energy and entropy of a system. The concept of a storage function originates in Willems’ theory of dissipative systems Willems (1972). In this context, reversible components as well as all EPHS without any irreversible components are said to be lossless, while (all systems with) irreversible components are said to be dissipative.
9 Example
This section compares a bond graph and an EPHS model of a DC shunt motor. A direct current (DC) motor has a fixed outer part, called stator, and a rotating inner part, called rotor. Both parts contain a coil of wire through which current flows. In the case of a shunt motor, these two inductors are connected in parallel (aka in shunt) to a source of electric energy. The working principle of such motors is based on a coupling of the electro-magnetic and the kinetic energy domains via the Lorentz force and the commutation of the rotor coil after every half turn. However, both aspects are not directly taken into account by the two models, which instead consider the coupling as being lumped into a gyrator element.
@@ 9 Example
We start with the bond graph model that is taken from Borutzky (2010) and redrawn in Figure 10. The half arrows are called bonds and each bond has two associated variables, called flow and effort. The name of the flow variable is written on the side of the arrow head, while the name of the effort variable is written on the opposite side. However, not all variables are explicitly named in the example. At any 00-junction the effort variables of all connected bonds are equal, while the flow variables balance, with their sign given by the bond direction. Dually, at $1$-junctions flows are equal and efforts balance. Specifically, at the 00-junction on the left the current from the $\mathrm{Se}$-element (effort source) with effort/voltage $E$ splits into the current $i_{s}$ through the stator coil and the current $i_{r}$ through the rotor coil. The inductance of the rotor coil is represented by the $\mathrm{I}$-element (generalized inductor) with inductance $L_{r}$, while an $\mathrm{I}$-element representing the stator coil inductance is omitted. The (steady-state) stator coil current $i_{s}$ is instead entirely determined by the $\mathrm{R}$-element (generalized resistor) at the top representing the coil’s electrical resistance $R_{s}$. The magnetic flux of the stator coil $b_{s}$ is instead obtained as the signal that results from multiplying the signal $i_{s}$, which is the shared effort variable at the upper $1$-junction, with the constant gain $L_{s}$. The bond graph hence combines energy-based modeling with signal-based modeling known from block diagrams. At the $1$-junction on the left, the source voltage $E$ is equal to the sum of the voltage over the rotor coil inductance $L_{r}$, the voltage $u_{R}$ over the rotor coil resistance $R_{r}$, and the induced (back EMF) voltage $u_{r}$. The latter is determined by the $\mathrm{MGY}$-element (modulated gyrator), which describes the coupling of the electrical and the mechanical parts. The gyrator ratio is the signal $b_{s}$ coming from the gain block. At the gyrator, the effort variable of one port is equal to the product of the gyrator ratio and the flow variable of the other port. On the mechanical side, the load torque $M_{\text{load}}\leq 0$ is modeled by the effort source on the right, while the rotor’s angular mass $J_{r}$ is represented by an $\mathrm{I}$-element. At the $1$-junction the sum of the generated torque $M$ and the load torque $M_{\text{load}}$ is equal to the sum of the torque accelerating the angular mass and the friction torque $M_{R}$. The latter is the product of the friction coefficient $d_{r}$ and the angular velocity $\omega$. Putting all this together, the bond graph gives the system of ordinary differential equations
@@
As already shown in 4.3, the EPHS model is specified as an interconnection of two subsystems named $\mathtt{stator}$ and $\mathtt{rotor}$. The two systems share an electric energy domain represented by the junction with connected ports named $\mathtt{q}$. The kinetic energy domain of the rotor is represented by the junction with connected ports named $\mathtt{p}$. Both energy domains are exposed such that the motor model can be integrated into another system. The conversion between electric and kinetic energy in the rotor depends on the state of the magnetic energy domain of the stator, which is represented by the junction with connected state ports named $\mathtt{b_{s}}$.
@@
The interconnection pattern of the stator model is shown in Figure 11. The box $\mathtt{em}$ is filled by a reversible component that represents the coupling of the electric energy domain with port $\mathtt{q}$ and the magnetic energy domain with port $\mathtt{b_{s}}$. The box $\mathtt{coil}$ is filled by a storage component that represents the inductance of the stator coil, while the box $\mathtt{emloss}$ is filled by an irreversible component that represents its electric resistance. The box $\mathtt{env}$ is filled by an environment component whose thermal energy domain with port $\mathtt{s}$ absorbs the dissipated electromagnetic energy.
The reversible component $(I_{\text{em}},\,\mathcal{D}_{\text{em}})$ is defined by its interface $I_{\text{em}}=(\{\mathtt{q},\,\mathtt{b_{s}}\},\,\tau_{\text{em}})$ with $\tau_{\text{em}}(\mathtt{q})=((\mathbb{R},\,\mathtt{charge}),\,\mathsf{p})$, $\tau_{\text{em}}(\mathtt{b_{s}})=((\mathbb{R},\,\mathtt{flux\_linkage}),\,\mathsf{p})$ and its Dirac structure $\mathcal{D}_{\text{em}}\subset\mathcal{P}_{I_{\text{em}}}$ given by
@@
The storage component $(I_{\text{coil}},\,E_{\text{coil}})$ is defined by its interface $I_{\text{coil}}=(\{\mathtt{b_{s}}\},\,\tau_{\text{coil}})$ with $\tau_{\text{coil}}(\mathtt{b_{s}})=((\mathbb{R},\,\mathtt{flux\_linkage}),\,\mathsf{p})$ and its energy function $E_{\text{coil}}:\mathcal{X}_{I_{\text{coil}}}\rightarrow\mathbb{R}$ with $\mathcal{X}_{I_{\mathtt{coil}}}=\mathbb{R}\ni\mathtt{b_{s}.x}=b_{s}$ given by
@@
where the parameter $L_{s}$ is the inductance. Since $H_{\mathtt{coil}}=E_{\mathtt{coil}}$, the effort variable $\mathtt{b_{s}.e}=\frac{b_{s}}{L_{s}}$ is the current through the coil.
The irreversible component $(I_{\text{emloss}},\,\mathcal{O}_{\text{emloss}})$ is defined by $I_{\text{emloss}}=(\{\mathtt{b_{s}},\,\mathtt{s}\},\,\tau_{\text{emloss}})$ with $\tau_{\text{emloss}}(\mathtt{b_{s}})=((\mathbb{R},\,\mathtt{flux\_linkage}),\,\mathsf{p})$, $\tau_{\text{emloss}}(\mathtt{s})=((\mathbb{R},\,\mathtt{entropy}),\,\mathsf{p})$ and its Onsager structure $\mathcal{O}_{\text{emloss}}\subset\mathcal{P}_{\text{emloss}}$ given by
$\begin{bmatrix}\mathtt{b_{s}.f}\\ \mathtt{s.f}\end{bmatrix}\>=\>\frac{1}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}\,R_{s}\,\begin{bmatrix}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}+}\mathtt{s.e}&-\mathtt{b_{s}.e}\\ -\mathtt{b_{s}.e}&\frac{{(\mathtt{b_{s}.e})}^{2}}{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}+}\mathtt{s.e}}\end{bmatrix}\,\begin{bmatrix}\mathtt{b_{s}.e}\\ \mathtt{s.e}\end{bmatrix}\>=\>\begin{bmatrix}R_{s}\,\mathtt{b_{s}.e}\\ -\frac{R_{s}\,{(\mathtt{b_{s}.e})}^{2}}{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}+}\mathtt{s.e}}\end{bmatrix}\,,$
where the parameter $R_{s}$ is the electric resistance of the coil.
@@
The interconnection pattern of the rotor model is shown in Figure 12. The components filling the boxes $\mathtt{coil}$, $\mathtt{em}$, $\mathtt{emloss}$, and $\mathtt{env}$ are defined as for the stator model, except for a different port name, namely $\mathtt{b}$ instead of $\mathtt{b_{s}}$, and different parameters, namely $L_{r}$ instead of $L_{s}$ and $R_{r}$ instead of $R_{s}$. The box $\mathtt{mk}$ is filled by a reversible component that describes the coupling of the magnetic energy domain with port $\mathtt{b}$ and the (rotational) kinetic energy domain with port $\mathtt{p}$. This coupling depends on the magnetic flux $\mathtt{b_{s}.x}$ that is generated by the stator coil. The box $\mathtt{mass}$ is filled by a storage component that represents the rotor’s angular mass and the box $\mathtt{mloss}$ is filled by an irreversible component that represents its mechanical friction.
The reversible component $(I_{\text{mk}},\mathcal{D}_{\text{mk}})$ is defined by $I_{\text{mk}}=(\{\mathtt{b},\,\mathtt{p},\,\mathtt{b_{s}}\},\,\tau_{\text{mk}})$ with $\tau_{\text{mk}}(\mathtt{b})=((\mathbb{R},\,\mathtt{flux\_linkage}),\,\mathsf{p})$, $\tau_{\text{mk}}(\mathtt{p})=((\mathbb{R},\,\mathtt{angular\_momentum}),\,\mathsf{p})$, $\tau_{\text{mk}}(\mathtt{b_{s}})=((\mathbb{R},\,\mathtt{flux\_linkage}),\,\mathsf{s})$ and its Dirac structure $\mathcal{D}_{\mathtt{mk}}\subset\mathcal{P}_{I_{\text{mk}}}$ given by
@@
The storage component $(I_{\text{mass}},\,E_{\text{mass}})$ is defined by $I_{\text{mass}}=(\{\mathtt{p}\},\,\tau_{\text{mass}})$ with $\tau_{\text{mass}}(\mathtt{p})=((\mathbb{R},\,\mathtt{angular\_momentum}),\,\mathsf{p})$ and $E_{\text{mass}}:\mathcal{X}_{I_{\text{mass}}}\rightarrow\mathbb{R}$ with $\mathcal{X}_{I_{\text{mass}}}=\mathbb{R}\ni\mathtt{p.x}=p$ given by
@@
The irreversible component $(I_{\text{mloss}},\mathcal{O}_{\text{mloss}})$ is defined by $I_{\text{mloss}}=(\{\mathtt{p},\,\mathtt{s}\},\,\tau_{\text{mloss}})$ with $\tau_{\text{mloss}}(\mathtt{p})=((\mathbb{R},\,\mathtt{angular\_momentum}),\,\mathsf{p})$, $\tau_{\text{mloss}}(\mathtt{s})=((\mathbb{R},\,\mathtt{entropy}),\,\mathsf{p})$ and its Onsager structure $\mathcal{O}_{\text{mloss}}\subset\mathcal{P}_{I_{\text{mloss}}}$ given by
$\begin{bmatrix}\mathtt{p.f}\\ \mathtt{s.f}\end{bmatrix}\>=\>\frac{1}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}\,d_{r}\,\begin{bmatrix}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}+}\mathtt{s.e}&-\mathtt{p.e}\\ -\mathtt{p.e}&\frac{{(\mathtt{p.e})}^{2}}{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}+}\mathtt{s.e}}\end{bmatrix}\,\begin{bmatrix}\mathtt{p.e}\\ \mathtt{s.e}\end{bmatrix}\>=\>\begin{bmatrix}d_{r}\,\mathtt{p.e}\\ -\frac{d_{r}\,{(\mathtt{p.e})}^{2}}{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}+}\mathtt{s.e}}\end{bmatrix}\,,$
where the parameter $d_{r}$ is the mechanical friction coefficient.
We write $(b_{s},\,b,\,p,\,s)\in\mathbb{R}^{4}$ for the state of the interconnected motor model, where we identify the entropy state variables $s$ of both environment components. In future work, the identification of environment components can be formalized by composing with a suitable relation. The EPHS model then yields the equations
@@
$\dot{s}\>$ $=\>\frac{1}{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}}\left(R_{s}\,i_{s}^{2}+R_{r}\,i_{r}^{2}+d_{r}\,\omega^{2}\right)$
$\mathtt{q.f}\>$ $=\>i_{s}+i_{r}\,,$
@@
$\omega\>$ $=\>\frac{1}{J_{r}}\,p\,.$

### v2

@@ 8 Thermodynamic consistency
Akin to metriplectic systems or the GENERIC, the structure of EPHS ensures thermodynamic consistency. To simplify things a bit, we focus on isolated systems.
@@
The dynamics of an isolated EPHS respect the first law and the second law. ∎
To prove this, one could first show that the semantics of any (isolated) composite system can be represented by some abstract system of evolution equations. Analogous to metriplectic systems or the GENERIC, one could then use the structural properties of this abstract system of equations to demonstrate thermodynamic consistency. The following proof, however, does not require this intermediate step.
@@
As the semantics associated to interconnection patterns is functorial, it suffices to show thermodynamic consistency for composite systems, where all subsystems are primitive. In terms of semantics, environment components are a special case of storage components. Any finite number of storage components can be combined into a single component by taking the named sum of their interfaces and by summing their energy functions. Similarly, any finite number of reversible/irreversible components can be combined into a single such component by taking the named sum of their interfaces and by taking the Cartesian product of the relations that define them. It hence suffices to show thermodynamic consistency for a composite system with one storage component, one reversible component, and one irreversible component.
@@
    First, we consider the storage component and focus specifically on the structure of its exergy function $H$, see 6.2. The following two facts follow immediately:
    1. (a)
      The energy $E$ must be conserved if the exergy $H$ and all quantities also present in the reference environment are conserved.
    2. (b)
      Entropy must be growing at a non-negative rate if the exergy $H$ is decreasing at a non-negative rate, while the energy $E$ and all quantities also present in the reference environment, except for entropy, are conserved.
@@
    Regarding the interconnection of the components based on 4.2, we know that at each junction the flow variables balance, while the effort variables are equal. We can conclude the following:
    1. (a)
      The interconnection is power-preserving, i.e. it conserves exergy.
    2. (b)
      The interconnection conserves the quantities associated with the ports because each flow variable represents the rate of change of the corresponding quantity, and the flow variables balance.
    3. (c)
      The interconnection conserves energy. This follows from the previous two points and point 1. (a)
@@
    Next, we consider the reversible component defined by a Dirac structure $\mathcal{D}$, see 6.7. Since $\mathcal{D}$ is a power-preserving relation, it conserves the exergy $H$. One of the conditions in 6.9 guarantees that $\mathcal{D}$ conserves all quantities also present in the reference environment, in particular entropy. Considering also point 1. (a), it follows that $\mathcal{D}$ must also conserve the energy $E$.
  4. 4.
    Finally, we consider the irreversible component defined by an Onsager structure $\mathcal{O}$, see 6.12. As a consequence of the non-negative definiteness property, $\mathcal{O}$ destroys exergy at a non-negative rate. One of the conditions in 6.14 guarantees that $\mathcal{O}$ conserves the energy $E$. Another condition guarantees that $\mathcal{O}$ conserves all quantities also present in the reference environment, except for entropy. Considering also point 1. (b), it follows that $\mathcal{O}$ must produce entropy at a non-negative rate.
The composite system must respect the first law, since the interconnection, the reversible component, and the irreversible component all conserve energy.
The composite system must respect the second law, since the interconnection and the reversible component conserve entropy, while the irreversible component produces entropy at a non-negative rate. ∎
The concept of a storage function was introduced in Willems’ theory of dissipative systems [49]. In this framework, EPHS without irreversible components are classified as lossless, whereas EPHS that include irreversible components are considered dissipative or passive. The additional conditions defining reversible and irreversible components in EPHS extend the notion of dissipativity, ensuring full thermodynamic consistency.
9 Example: shunt motor
This section compares a bond graph and an EPHS model of a direct current (DC) shunt motor. A DC motor has a fixed outer part, called stator, and a rotating inner part, called rotor. Both parts contain a coil of wire through which current flows. In the case of a shunt motor, these two inductors are connected in parallel to a source of electric energy. The working principle of such motors is based on a coupling of the electromagnetic and the kinetic energy domains via the Lorentz force and the commutation of the rotor coil after every half turn. However, both aspects are not directly resolved by the considered models. Instead, the coupling is considered to be lumped into a gyrator element.
@@ 9 Example: shunt motor
We start with the bond graph model that is taken from [8] and reproduced, with slightly different naming of variables, in Figure 10. The half arrows are called bonds and each bond has two associated variables, called flow and effort. The name of the flow variable is written on the side of the arrow head, while the name of the effort variable is written on the opposite side. However, not all variables are explicitly named in the example. At any 00-junction, the effort variables of all connected bonds are equal, while the flow variables balance, with their sign given by the bond direction. Dually, at $1$-junctions, flows are equal and efforts balance. Specifically, at the 00-junction on the left, the current from the $\mathrm{Se}$-element (effort source) with effort/voltage $E$ splits into the current $i_{s}$ through the stator coil and the current $i_{r}$ through the rotor coil. The inductance of the rotor coil is represented by the $\mathrm{I}$-element (generalized inductor) with inductance $L_{r}$, while an $\mathrm{I}$-element representing the stator coil inductance is omitted. The (steady-state) stator coil current $i_{s}$ is instead entirely determined by the $\mathrm{R}$-element (generalized resistor) at the top representing the coil’s electrical resistance $R_{s}$. The magnetic flux of the stator coil $b_{s}$ is then obtained as the signal that results from multiplying the signal $i_{s}$, which is the shared effort variable at the upper $1$-junction, with the constant gain $L_{s}$. The bond graph hence combines energy-based modeling with signal-based modeling based on block diagrams. At the $1$-junction on the left, the source voltage $E$ is equal to the sum of the voltage over the rotor coil inductance $L_{r}$, the voltage $u_{R}$ over the rotor coil resistance $R_{r}$, and the induced (back EMF) voltage $u_{r}$. The latter is determined by the $\mathrm{MGY}$-element (modulated gyrator), which describes the coupling of the electrical and the mechanical parts. The gyrator ratio is the signal $b_{s}$ coming from the gain block. At the gyrator, the effort variable of one port is equal to the product of the gyrator ratio and the flow variable of the other port. On the mechanical side, the load torque $M_{\text{load}}\leq 0$ is modeled by the effort source on the right, while the rotor’s angular mass $J_{r}$ is represented by an $\mathrm{I}$-element. At the $1$-junction the sum of the generated torque $M$ and the load torque $M_{\text{load}}$ is equal to the sum of the torque accelerating the angular mass and the friction torque $M_{R}$. The latter is the product of the friction coefficient $d_{r}$ and the angular velocity $\omega$. Putting all this together, the bond graph gives the system of ordinary differential equations
@@
As already shown in 4.3, the EPHS model is specified as an interconnection of two subsystems named $\mathtt{stator}$ and $\mathtt{rotor}$. The two systems have a common electric energy domain represented by the junction with connected ports named $\mathtt{q}$. The kinetic energy domain of the rotor is represented by the junction with connected ports named $\mathtt{p}$. Both energy domains are exposed such that the motor model can be integrated into another system. The conversion between electric and kinetic energy in the rotor depends on the state of the magnetic energy domain of the stator, which is represented by the junction with connected state ports named $\mathtt{b_{s}}$.
@@
The interconnection pattern of the stator model is shown in Figure 11. Box $\mathtt{emc}$ is filled by a reversible component that represents the coupling of the electric energy domain associated with port $\mathtt{q}$ and the magnetic energy domain associated with port $\mathtt{b_{s}}$. Box $\mathtt{coil}$ is filled by a storage component that represents the inductance of the stator coil. Further, box $\mathtt{res}$ is filled by an irreversible component that represents the resistance of the coil. Finally, box $\mathtt{env}$ is filled by an environment component whose thermal energy domain associated with port $\mathtt{s}$ absorbs the dissipated electromagnetic energy.
The reversible component $(I_{\text{emc}},\,\mathcal{D}_{\text{emc}})$ is defined by its interface
$I_{\text{emc}}=(\{\mathtt{q},\,\mathtt{b_{s}}\},\,\tau_{\text{emc}})$ with
$\tau_{\text{emc}}(\mathtt{q})\>$ $=\>((\mathbb{R},\,\mathtt{charge}),\,\mathsf{p})$
$\tau_{\text{emc}}(\mathtt{b_{s}})\>$ $=\>((\mathbb{R},\,\mathtt{flux\_linkage}),\,\mathsf{p})$
and its Dirac structure $\mathcal{D}_{\text{emc}}\subset\mathcal{P}_{I_{\text{emc}}}$ given by
@@
The storage component $(I_{\text{coil}},\,E_{\text{coil}})$ is defined by its interface $I_{\text{coil}}=(\{\mathtt{b_{s}}\},\,\tau_{\text{coil}})$ with $\tau_{\text{coil}}(\mathtt{b_{s}})=((\mathbb{R},\,\mathtt{flux\_linkage}),\,\mathsf{p})$ and its energy function $E_{\text{coil}}:\mathcal{X}_{I_{\text{coil}}}\rightarrow\mathbb{R}$ with $\mathcal{X}_{I_{\mathtt{coil}}}=\mathbb{R}\ni\mathtt{b_{s}.x}=b_{s}$. The function is given by
@@
where the parameter $L_{s}$ is the inductance. The effort variable $\mathtt{b_{s}.e}=\frac{b_{s}}{L_{s}}$ is the current through the coil.
The irreversible component $(I_{\text{res}},\,\mathcal{O}_{\text{res}})$ is defined by its interface $I_{\text{res}}=(\{\mathtt{b_{s}},\,\mathtt{s}\},\,\tau_{\text{res}})$ with
$\tau_{\text{res}}(\mathtt{b_{s}})\>$ $=\>((\mathbb{R},\,\mathtt{flux\_linkage}),\,\mathsf{p})$
$\tau_{\text{res}}(\mathtt{s})\>$ $=\>((\mathbb{R},\,\mathtt{entropy}),\,\mathsf{p})$
and its Onsager structure $\mathcal{O}_{\text{res}}\subset\mathcal{P}_{\text{res}}$ given by
$\begin{bmatrix}\mathtt{b_{s}.f}\\ \mathtt{s.f}\end{bmatrix}\>=\>\frac{1}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}\,R_{s}\,\begin{bmatrix}\theta&-i\\ -i&\frac{i^{2}}{\theta}\end{bmatrix}\,\begin{bmatrix}\mathtt{b_{s}.e}\\ \mathtt{s.e}\end{bmatrix}\>=\>\begin{bmatrix}R_{s}\,i\\ -\frac{1}{\theta}\,R_{s}\,i^{2}\end{bmatrix}\,,$
where the parameter $R_{s}$ is the resistance of the coil, $i=\mathtt{b_{s}.e}$ is the current through the coil, and $\theta={\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}+}\mathtt{s.e}$ is the absolute temperature at which heat is dissipated into the thermal energy domain.
@@
The interconnection pattern of the rotor model is shown in Figure 12. The components filling the boxes $\mathtt{coil}$, $\mathtt{emc}$, $\mathtt{res}$ and $\mathtt{env}$ are defined as for the stator model, except for a different port name, namely $\mathtt{b}$ instead of $\mathtt{b_{s}}$, and different parameters, namely $L_{r}$ instead of $L_{s}$ and $R_{r}$ instead of $R_{s}$. Further, box $\mathtt{mkc}$ is filled by a reversible component that describes the coupling of the magnetic energy domain associated with port $\mathtt{b}$ and the (rotational) kinetic energy domain associated with port $\mathtt{p}$. This coupling depends on the magnetic flux $\mathtt{b_{s}.x}$ that is generated by the stator coil. Box $\mathtt{mass}$ is filled by a storage component that represents the rotor’s angular mass and box $\mathtt{mf}$ is filled by an irreversible component that represents its mechanical friction.
The reversible component $(I_{\text{mkc}},\mathcal{D}_{\text{mkc}})$ is defined by its interface
$I_{\text{mkc}}=(\{\mathtt{b},\,\mathtt{p},\,\mathtt{b_{s}}\},\,\tau_{\text{mkc}})$ with
$\tau_{\text{mkc}}(\mathtt{b})\>$ $=\>((\mathbb{R},\,\mathtt{flux\_linkage}),\,\mathsf{p})$
$\tau_{\text{mkc}}(\mathtt{p})\>$ $=\>((\mathbb{R},\,\mathtt{angular\_momentum}),\,\mathsf{p})$
$\tau_{\text{mkc}}(\mathtt{b_{s}})\>$ $=\>((\mathbb{R},\,\mathtt{flux\_linkage}),\,\mathsf{s})$
and its Dirac structure $\mathcal{D}_{\mathtt{mkc}}\subset\mathcal{P}_{I_{\text{mkc}}}$ given by
@@
The storage component $(I_{\text{mass}},\,E_{\text{mass}})$ is defined by its interface $I_{\text{mass}}=(\{\mathtt{p}\},\,\tau_{\text{mass}})$ with $\tau_{\text{mass}}(\mathtt{p})=((\mathbb{R},\,\mathtt{angular\_momentum}),\,\mathsf{p})$ and its energy function
$E_{\text{mass}}:\mathcal{X}_{I_{\text{mass}}}\rightarrow\mathbb{R}$ with $\mathcal{X}_{I_{\text{mass}}}=\mathbb{R}\ni\mathtt{p.x}=p$. The function is given by
@@
The irreversible component $(I_{\text{mf}},\mathcal{O}_{\text{mf}})$ is defined by its interface $I_{\text{mf}}=(\{\mathtt{p},\,\mathtt{s}\},\,\tau_{\text{mf}})$ with
$\tau_{\text{mf}}(\mathtt{p})\>$ $=\>((\mathbb{R},\,\mathtt{angular\_momentum}),\,\mathsf{p})$
$\tau_{\text{mf}}(\mathtt{s})\>$ $=\>((\mathbb{R},\,\mathtt{entropy}),\,\mathsf{p})$
and its Onsager structure $\mathcal{O}_{\text{mf}}\subset\mathcal{P}_{I_{\text{mf}}}$ given by
$\begin{bmatrix}\mathtt{p.f}\\ \mathtt{s.f}\end{bmatrix}\>=\>\frac{1}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}\,d_{r}\,\begin{bmatrix}\theta&-\upsilon\\ -\upsilon&\frac{\upsilon^{2}}{\theta}\end{bmatrix}\,\begin{bmatrix}\mathtt{p.e}\\ \mathtt{s.e}\end{bmatrix}\>=\>\begin{bmatrix}d_{r}\,\upsilon\\ -\frac{d_{r}\,\upsilon^{2}}{\theta}\end{bmatrix}\,,$
where the parameter $d_{r}$ is the mechanical friction coefficient, $\upsilon=\mathtt{p.e}$ is the angular velocity, and $\theta={\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}+}\mathtt{s.e}$ is the absolute temperature at which heat is dissipated into the thermal energy domain.
We write $(b_{s},\,b,\,p,\,s)\in\mathbb{R}^{4}$ for the state of the interconnected motor model, where we identify the entropy state variables $s$ of both environment components. In a future work, the identification of environment components can be formalized by composing with a suitable relation. The EPHS model then yields the equations
@@
$\dot{s}\>$ $=\>\frac{1}{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}}\left(R_{s}\,i_{s}^{2}+R_{r}\,i_{r}^{2}+d_{r}\,\omega^{2}\right)$
$\mathtt{q.f}\>$ $=\>i_{s}+i_{r}$
$\mathtt{p.e}\>$ $=\>\omega\,,$
@@
$\omega\>$ $=\>\frac{1}{J_{r}}\,p\,,$
which are similar to the equations for the bond graph model, except that (1) the source of electrical energy and the mechanical load are regarded as external dynamical systems, rather than internal constants, and (2) no quasi-static assumption is hence made for the magnetic field of the stator.